OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. Squareroot(16y^3) - Squareroot(y3)Please explain how! :(

OpenStudy (luigi0210):

\[\sqrt{16y^3}-\sqrt{y^3}\] This?

OpenStudy (anonymous):

oh wait yes! sorry i wrote the question wrong xD

OpenStudy (luigi0210):

Well just break it down to square roots that you know of: \[\sqrt{16y^2*y}-\sqrt{y^2*y}\]

OpenStudy (anonymous):

\[4\sqrt{y^2*y} - \sqrt{y^2*y} \] I'm sorry I'm kind of slow, I'm new here so xD Is this how you do it?

OpenStudy (anonymous):

The answer is \[3y^2 \sqrt{y}\] right?

OpenStudy (luigi0210):

How did the 4 turn into a 3? :P

OpenStudy (anonymous):

Oh oops, brain fart. Like i assume the answer is four, but in the multiple choices said it doesn't have the answer '4' So I'm a bit confused hahaha.. I'm probably wrong with the 4

OpenStudy (anonymous):

Is it \[4y \sqrt{y} \] >_<?

OpenStudy (luigi0210):

Sorry about that.. tech difficulties..